if two equal charges , each of 1c,were separated in air by a distance of 1km,what would be the force between them?

1
Expert's answer
2022-06-10T08:28:34-0400
"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9\frac{\rm N\cdot m^2}{\rm C^2}*\frac{1\: \rm C*1\: C}{(10^3\:\rm m)^2}=9000\:\rm N"